### v3

# Blending 3D Geometry and Machine Learning for Multi-View Stereopsis

Abstract
Traditional multi-view stereo (MVS) methods primarily depend on photometric and geometric consistency constraints. In contrast, modern learning-based algorithms often rely on the plane sweep algorithm to infer 3D geometry, applying explicit geometric consistency (GC) checks only as a post-processing step, with no impact on the learning process itself. In this work, we introduce GC-MVSNet++, a novel approach that actively enforces geometric consistency of reference view depth maps across multiple source views (multi-view) and at various scales (multi-scale) during the learning phase (see Fig. 1). This integrated GC check significantly accelerates the learning process by directly penalizing geometrically inconsistent pixels, effectively halving the number of training iterations compared to other MVS methods. Furthermore, we introduce a densely connected cost regularization network with two distinct block designs—simple and feature-dense—optimized to harness dense feature connections for enhanced regularization. Extensive experiments demonstrate that our approach achieves a new state-of-the-art on the BlendedMVS dataset, and competitive performance on the DTU and Tanks and Temples benchmark. To our knowledge, GC-MVSNet++ is among the few approaches that enforce supervised geometric consistency across multiple views and at multiple scales during training. Our code is available at https://github.com/vkvats/GC-MVSNet-PlusPlus
1 Introduction
Traditional multi-view stereo (MVS) methods, such as Gipuma [1], Furu [2], COLMAP [3], and Tola [4], primarily focused on addressing photometric and geometric consistency constraints across multiple views. In contrast, recent machine learning-based MVS approaches [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have used deep networks to extract feature maps and combine them into 3D cost volumes, capturing subtle similarities between features. These advances—ranging from multi-level feature extraction [5, 10] and attention-driven feature matching [5, 19] to refined cost-volume creation [5, 6, 8, 19] and innovative loss formulations [5, 6, 13]—have significantly improved the fidelity of depth estimates and point cloud reconstructions.
However, unlike the traditional approaches, these machine learning-based techniques typically model geometric consistency in only an indirect way. Typically, consistency checks are reserved for post-processing [5, 8, 10, 16], filtering depth maps only after inference is complete. This means the rich, multi-view geometric constraints are not explicitly modeled during learning, leaving the network to uncover these relationships on its own, often in subtle and less direct ways.
In this paper, we demonstrate that incorporating explicit multi-view geometric cues through geometric consistency checks across multiple source views during training (see Fig. 1) significantly enhances accuracy while reducing the number of training iterations. To further improve the MVS pipeline, we present a novel densely connected cost regularization network with two distinct block architectures, specifically designed to fully exploit dense feature connections during the cost regularization process. Additionally, we enhance the feature extraction network with weight-standardized deformable convolution, ensuring improved feature extraction.
These innovations lead to the development of GC-MVSNet++, a multi-stage model that learns geometric cues across three scales. At each scale, we integrate a multi-view geometric consistency module that performs geometric consistency checks on reference view depth estimates across multiple source views, generating a per-pixel penalty. This penalty is then combined with the per-pixel depth error, calculated using cross-entropy loss at each stage, to form the final loss function, guiding the model toward more precise reconstructions.
Figure 1: Our multi-view, multi-scale geometric consistency process. It explicitly models the geometric consistency of the estimated depth map across multiple source views during training at coarse, intermediate, and refine stages. At each stage, the ref. depth estimate map undergo forward-backward-reprojection (see Alg. 2) across NN src. views to generate the GC-penalty map to penalize per-pixel depth error. This approach enables the model to learn geometric constraints more quickly and accurately, leading to improved reconstruction quality during inference.
This formulation of the loss function offers richer geometric cues that significantly expedite model learning. Our extensive experiments reveal that GC-MVSNet++ nearly halves the number of training iterations compared to other contemporary models [5, 6, 10, 13, 15]. Our approach not only sets a new benchmark for accuracy on the BlendedMVS [20] dataset, but also secures the second position on the DTU [21] and Tanks and Temples [22] advanced benchmark. GC-MVSNet++ is novel in its use of multi-view, multi-scale geometric consistency checks during training with a novel cost regularization network. Extensive ablation experiments further underscore the efficacy of our proposed method. To summarize:
  • $\bullet$
    We introduce an innovative multi-view, multi-scale geometric consistency module that fosters geometric coherence throughout the learning process.
  • $\bullet$
    Our method reduces the training iterations by almost 50% compared to other models, thanks to its explicit geometric cues.
  • $\bullet$
    The versatility of GC-module allows it to be seamlessly integrated into various MVS pipelines to boost geometric consistency during training.
  • $\bullet$
    Additionally, the GC module serves as a pre-processing tool to eliminate geometrically inconsistent pixels from the ground truth data.
  • $\bullet$
    We also present a novel cost regularization network featuring two distinct block designs that leverage dense connections for cost regularization.
A preliminary conference version of this work appeared in [19]. In this updated version, we introduce an innovative cost regularization network (Sec. 3.3) and employ depth filtering (Sec. 3.7) to enhance GC-MVSNet and achieve better experimental results. We also offer additional architectural details and perform more extensive experimental evaluations.
2 Related Work
Furukawa and Ponce [2] propose a taxonomy that classifies MVS methods into four primary scene representations: volumetric fields [23, 24, 25, 26], point clouds [9, 27], 3D meshes [28], and depth maps [1, 3, 5, 6, 10, 12, 14, 15, 16, 29, 18, 30]. Depth map-based methods can be further divided into traditional techniques that rely on feature detection and geometric constraint solving [1, 2, 3, 29], and learning-based approaches [5, 6, 10, 12, 14, 15, 16], which have gained significant popularity in recent years.
Among the learning-based techniques, MVSNet [15] presents a single-stage MVS pipeline that encodes camera parameters through differential homography to construct 3D cost volumes. However, it demands substantial memory and computational resources due to its use of 3D U-Nets [31] for cost volume regularization. To address this challenge, subsequent research has pursued two primary strategies: the adoption of recurrent neural networks (RNNs) [8, 13, 32, 33] and the development of coarse-to-fine multi-stage methods [5, 6, 10, 12, 14].
Figure 2: The GC-MVSNet++ architecture. It incorporates the GC module at the end of each stage. This module utilizes the estimated reference view depth, $M$ source view ground truths, and their associated camera parameters to conduct a multi-view geometric consistency check. It generates a per-pixel penalty ($\xi_{p}$) for reference view, which is then element-wise multiplied with per-pixel depth error ($\xi_{d}$) to compute the stage loss $L_{i}$. $\xi_{d}$ is calculated using cross-entropy loss. Figs. 4 and 5 provide a detailed view of the Cost regularization network and feature extraction network, respectively.
Coarse-to-fine multi-stage methods [5, 6, 7, 8, 9, 10, 11, 12, 13] have markedly enhanced the quality of depth estimates and point cloud reconstructions. These methods start by predicting a low-resolution (coarse) depth map and then refine it progressively. For instance, CasMVSNet [10] extends the single-stage MVSNet [15] into a multi-stage framework, while TransMVSNet [5] improves performance over CasMVSNet through advanced feature matching techniques. UniMVSNet [6] further builds on CasMVSNet by employing a unified loss formulation to enhance accuracy. CVP-MVSNet [11] constructs a cost volume pyramid in a coarse-to-fine manner, and UCS-Net [10] introduces an adaptive thin volume module that optimally partitions the local depth range with fewer hypothesis planes. Additionally, TransMVSNet [5] incorporates transformer-based feature matching [34, 35] to improve feature similarity, while UniMVSNet [6] integrates the benefits of both regression and classification approaches through a unified focal loss within its multi-stage framework.
While these methods enhance multi-stage MVS pipelines by refining specific components, they generally lack explicit integration of multi-view geometric cues during the learning process. Consequently, these models depend on the limited geometric information from multiple source views and the cost function formulation during training. Xu and Tao [14] address this limitation with a multi-scale geometric consistency-guided MVS approach that uses multi-hypothesis joint view selection to leverage structured region information for improved candidate hypothesis sampling. They argue that upsampled depth maps from source images can impose geometric constraints on estimates and utilize reprojection error [3, 36, 37] to assess consistency. Unsupervised methods [38, 39] use cross-view consistency to learn geometric features. In contrast, our method employs forward-backward reprojection across multiple source views to directly evaluate and enforce geometric consistency in depth estimates, generating per-pixel penalties for geometrically inconsistent pixels.
Cost volume regularization: Cost volume regularization networks enhance raw cost volumes by converting them into regularized (smooth) volumes prior to depth estimation. Early approaches utilized the 3D-UNet architecture for similar regularization in stereo methods [40, 41, 42]. These techniques typically employed variations of a basic encoder-decoder network. For instance, PSMNet [41] introduced a three-tier Hourglass architecture, while GANet [42] incorporated a cost aggregation block featuring semi-global and local guided modules. Additionally, Guo et al. [43] combined correlation and cost volume methods to leverage their respective strengths.
Building on its success in stereo and MVS tasks [40, 44, 45], MVSNet applies a multi-scale 3D-UNet for cost volume regularization. R-MVSNet [8] enhances this with a recurrent GRU model to refine the cost volumes. CasMVSNet [10] uses separate networks for each stage, guiding each subsequent stage with the depth estimate from the previous one. PointMVSNet [9] further extends this approach to four stages. However, many recent methods [5, 6, 10, 13, 19, 46] do not significantly innovate in the architecture of cost regularization networks, typically employing simple encoder-decoder structures with 3D convolutional layers. In contrast, we propose an advanced densely connected U-Net architecture featuring two distinct block designs for the encoding and decoding phases.
Inputs: $D_{0},c_{0},D^{gt}_{i},c^{gt}_{i},D_{pixel},D_{depth}$
Output: per_pixel_penaltyper\_pixel\_penalty
Require $M\geq N$
$mask\_sum\leftarrow 0$
$D\leftarrow{D_{1}^{gt},...D_{M}^{gt}}$
$c\leftarrow{c_{1}^{gt},...c_{M}^{gt}}$
for $D_{i}^{gt},c_{i}^{gt}$ in $zip(D,c)$ do
  $D^{\prime\prime}_{P^{\prime\prime}_{0}},P^{\prime\prime}_{0}\leftarrow FBR(D_{0},c_{0},D_{i}^{gt},c_{i}^{gt})$ {Alg. 2}
[…]
  $PDE_{mask}\leftarrow{PDE>D_{pixel}}$
  $RDD_{mask}\leftarrow{RDD>D_{depth}}$
  $mask\leftarrow PDE_{mask}\lor RDD_{mask}$
  if $mask>0$ then
   $mask\leftarrow 1$
  else
   $mask\leftarrow 0$
  end if
  $mask\_sum\leftarrow mask\_sum+mask$
[…]
Algorithm 1 Geometric Consistency Check Algorithm
Inputs: $D_{0},c_{0},D_{i}^{gt},c_{i}^{gt}$
Output: $D^{\prime\prime}_{P^{\prime\prime}_{0}},P^{\prime\prime}_{0}$
$K_{R},E_{R}\leftarrow c_{0}$; $K_{S},E_{S}\leftarrow c_{i}^{gt}$
$D_{(R\rightarrow S)}\leftarrow K_{S}\cdot E_{S}\cdot E_{R}^{-1}\cdot K_{R}^{-1}\cdot D_{0}$ {Project}
$X_{D_{(R\rightarrow S)}},Y_{D_{(R\rightarrow S)}}\leftarrow D_{(R\rightarrow S)}$
$D_{S_{remap}}\leftarrow REMAP(D_{i}^{gt},X_{D_{(R\rightarrow S)}},Y_{D_{(R\rightarrow S)}})$ {Remap}
$D^{\prime\prime}_{P^{\prime\prime}_{0}}\leftarrow K_{R}\cdot E_{R}\cdot E_{S}^{-1}\cdot K_{S}^{-1}\cdot D_{S_{remap}}$ {Back project}
$P^{\prime\prime}_{0}\leftarrow(X_{D^{\prime\prime}_{P^{\prime\prime}_{0}}},Y_{D^{\prime\prime}_{P^{\prime\prime}_{0}}})$
Algorithm 2 Forward Backward Reprojection (FBR)
3 Methodology
MVS methods take $N$ views as input, including a reference image $I_{0}\in\mathbb{R}^{H\times W\times 3}$ and its paired (N1N-1) source view images $I^{N-1}_{i=1}$, along with the corresponding camera parameters $c_{0},...,c_{N}$, and then to estimate the reference view depth map ($D_{0}$) as the output.
3.1 Network Overview
The architecture of our approach Geometric Consistency MVSNet++, abbreviated as GC-MVSNet++, is shown in Fig. 2. We use a deformable convolution-based [47] feature pyramid network (FPN) [48] architecture (Sec. 3.4) to extract features from input images in a coarse-to-fine manner in three stages. At only the coarse stage, we apply a Global Feature Aggregation (GFA) module with linear attention [35, 5] to leverage global context information within and between reference and source view features. At each stage, we build a correlation-based cost volume of shape $H^{\prime}\times W^{\prime}\times D^{\prime}_{h}\times 1$ using feature maps of shape $N\times H^{\prime}\times W^{\prime}\times C$, where $H^{\prime}$, $W^{\prime}$, and $C$ denote the height, width, and number of channels of a given stage, and $D^{\prime}_{h}$ is the number of depth hypotheses at the corresponding stage. The cost volume is regularized with the proposed dense-connected CostRegNet. We use a winner-takes-all strategy to estimate the depth map $D_{0}$ at each stage.
We employ the GC module at each stage. It checks the geometric consistency of each pixel in $D_{0}$ across $M$ source views and generates $\xi_{p}$ (Sec. 3.2), a pixel-wise factor that is multiplied with the per-pixel depth error ($\xi_{d}$), calculated using a cross-entropy function. It penalizes each pixel in $D_{0}$ for its inconsistency across $M$ source views to accelerate geometric cues learning during training. TransMVSNet [5] (without feature matching transformer and adaptive receptive field module) trained with cross-entropy loss is our baseline (GC-MVS-base), see Table 7.
Figure 3: The final $\xi_{p}$ is the outcome of elementwise multiplication ($\odot$) of initial $\xi_{p}$ and reference view mask. It restricts the penalties within the reference view mask.
3.2 Multi-View Geometric Consistency Module
Figure 4: Expanded view of the proposed Cost Regularization Network (CostRegNet) from Fig. 2. (a) shows the architecture of the dense CostRegNet with three types of blocks – dense, feature-dense, and transition blocks. (b) shows the internal architecture of the transition block, its components include a weight-standardized 3D convolution layer followed by a 3D dropout and 3D max-pooling layers. (c) and (d) show a single unit inside the feature dense and dense units. (e) shows a four-layer dense block and (f) shows a three-layer feature dense block.
GC-MVSNet++ estimates reference depth maps at three stages with different resolutions. At each stage, the GC module takes $D_{0}$, $M$ source view ground truths $D^{gt}_{1},...D^{gt}_{M}$, and their camera parameters $c_{0},...c_{M}$ as input (see Alg. 1). The GC module is then initialized with a geometric inconsistency mask sum (or mask_summask\_sum) of zero at each stage. This mask sum accumulates the inconsistency of each pixel across the $M$ source views. For each source view, the GC module performs forward-backward reprojection of $D_{0}$ to generate the penalty and then adds it to the mask sum.
Forward-backward reprojection (FBR), as shown in Alg. 2, is a crucial three-step process. First, we project each pixel $P_{0}$ of $D_{0}$ to its $i^{th}$ neighboring source view using intrinsic ($K_{R},K_{S}$) and extrinsic ($E_{R},E_{S}$) camera parameters to obtain corresponding pixel $P^{\prime}_{i}$, and denote the corresponding depth map as $D_{(R\rightarrow S)}$. Second, we similarly remap $D_{i}^{gt}$ to obtain $D_{S_{remap}}$. Finally, we back project $D_{S_{remap}}$ to the reference view using intrinsic and extrinsic camera parameters to obtain $D^{\prime\prime}_{P^{\prime\prime}_{0}}$ (see Alg. 2). $D_{0}$ and $D^{\prime\prime}_{P^{\prime\prime}_{0}}$ represent the depth values of pixels $P_{0}$ and $P^{\prime\prime}_{0}$ [49]. With $P^{\prime\prime}_{0}$ and $D^{\prime\prime}_{P^{\prime\prime}_{0}}$, we calculate the pixel displacement error (PDE) and relative depth difference (RDD). PDE is the $L_{2}$ norm between $P_{0}$ and $P^{\prime\prime}_{0}$ and RDD is the absolute value difference between $D^{\prime\prime}_{P^{\prime\prime}_{0}}$ and $D_{0}$ relative to $D_{0}$ as shown in Alg. 1.
For each stage, we generate two binary masks of inconsistent pixels, $PDE_{mask}$ and $RDD_{mask}$, by applying thresholds $D_{pixel}$ and $D_{depth}$, and then take a logical-OR of the two to produce a single mask of inconsistent pixels. These inconsistent pixels are assigned a value $1$ and all other pixels, including the consistent and the out-of-scope pixels, are assigned 0 to form a penalty mask. This penalty mask is then added to the mask sum (Alg. 1), which accumulates the penalty mask for each of the $M$ source views to generate a final mask sum with values $\in[0,M]$. Each pixel value indicates the number of inconsistencies of the pixel across the $M$ source views.
From this mask sum, we then generate the inconsistency penalty $\xi_{p}$ for each pixel. Our initial approach generated $\xi_{p}$ by dividing the mask sum by $M$ to normalize within the $[0,1]$. However, we found that using $\xi_{p}$ itself for element-wise multiplication reduces the contribution of perfectly consistent (zero inconsistency) pixels to zero, preventing further improvement of such pixels. To avoid this, we add $1$ so that elements of $\xi_{p}\in[1,2]$. A reference view binary mask is applied on initial $\xi_{p}$ to generate the final $\xi_{p}$, as shown in Fig. 3.
Occlusion and its impact: In multi-view stereo, occluded pixels naturally occur as 3D points often remain invisible from certain views. These hidden pixels significantly affect geometric constraints, leading to the penalization of reference view pixels corresponding to occluded 3D points as inconsistent. To prevent these occluded pixels from disproportionately impacting geometric consistency losses, careful management is essential.
Although some methods explicitly model occlusion [50, 51], our approach is inherently resilient to occlusion due to three key factors. First, we select the closest $M$ source views, as defined in MVSNet [15], to reduce the occurrence of occluded pixels across different views. During the FBR process, we remap $D_{i}^{gt}$ to generate $D_{S_{remap}}$ and perform back projection as detailed in Alg. 2. This remapping and back projection effectively manage severe occlusions (refer to Appendix A in Supplemental Material). Lastly, we use a binary mask on $\xi_{p}$, as illustrated in Fig. 3, to confine penalties to valid reference view pixels only. These combined strategies enable us to address the challenges posed by occluded pixels and prevent the explosion of loss values.
Figure 5: Expanded view of the Deformable Feature Extraction Network (FEN) and its components. (a) shows the overall architecture of the FEN. (b) and (c) shows the internal design of the deformable-conv2d and stage-out block. (d) shows the two distinct methods of global feature aggregation for reference and source image features.
3.3 Cost Regularization Network (CostRegNet)
The raw cost volume derived from reference and source image features often contains significant noise, originating from occlusions, feature mismatches, and non-Lambertian surfaces. This noise can hinder the accuracy of depth estimation, necessitating regularization to achieve smoother results. To address this, we introduce a novel densely connected [52] cost regularization network, termed dense-CostRegNet, designed specifically for dense prediction tasks.
Dense-CostRegNet leverages the DenseNet [52] architecture, known for its dense block connections. DenseNet was developed as an enhancement over ResNet [53], focusing on improved image recognition through its dense connectivity within blocks. Drawing inspiration from its success in image recognition, we have adapted this concept into a U-Net [31] style encoder-decoder framework. Our design features distinct encoder and decoder blocks tailored for depth estimation challenges, as illustrated in Fig. 4 (a).
In the U-Net architecture, the encoding phase employs a simple dense block, closely adhering to the original DenseNet design. In contrast, the bottleneck and decoder stages utilize a feature-dense block. This feature-dense block emphasizes generating new features while capitalizing on the benefits of dense feature connections within the block, as depicted in Fig. 4 (c) and (f). We believe that the dense connections within these blocks are crucial for enhancing the smoothness of the cost volume, which leads to more accurate depth estimates. Below, we provide a detailed explanation of each component of the dense-CostRegNet, beginning with an overview of dense connectivity and its advantages.
3.3.1 Dense Connectivity
Let us assume $H_{k}(.)$ denotes the feed-forward 2D convolution operation in a traditional 2D convolutional neural network (CNN), which feeds the output feature map $x_{k-1}$ of the $(k-1)^{th}$ layer to the input of the next layer, $x_{k}=H_{k-1}(x_{k-1})$. To bypass the non-linear transformation, ResNet [53] additionally adds skip-connections, $x_{k}=H_{k-1}(x_{k-1})+x_{k-1}$. DenseNet [52] further improves the information flow between layers by adding a denser connectivity pattern with direct connections from any layer to all its subsequent layers. More precisely, the $k^{th}$ layer receives the concatenation of feature maps $x_{0},x_{1},...x_{k-1}$] from all its preceding layers,
$x_{k}=H([x_{0},x_{1},...x_{k-1}]),$ (1)
$H_{k}(.)$ is a composite of sequential layer operations. The output feature dimension of the $k^{th}$ layer depends on $g$ (a fixed growth rate parameter). This can be easily extended to 3D convolutional layers.
3.3.2 Encoder Dense Blocks
The encoder uses a 3D version of the simple formulation of a dense block as discussed in Eq. 1. It reduces the feature resolution to $\frac{1}{8}$ of the input to reach the bottleneck. Each stage of the encoder (with fixed feature resolution) has only one dense block but uses a different number of dense-block units. Fig. 4 (c) shows the design of a single dense-block unit with weight-standardized 3D convolution, followed by a group normalization layer, ReLU non-linearity, and a 3D dropout layer. The learned features ($x^{\prime}$) are concatenated with the input feature ($x_{in}$) for the next dense-block unit. Fig. 4 (e) shows a 4-layer dense block. The number of layers (dense units) inside each block is shown as a superscript in the U-Net diagram (Fig. 4(a)).
3.3.3 Decoder Dense Blocks
The bottleneck and the decoder parts of the U-Net use the same feature-dense block. Fig. 4 (c) shows the internal design of a unit. It includes a weight-standardized 3D convolution followed by a GroupNorm [54] and ReLU non-linearity. The learned $g$ (dense block growth rate) new features $x^{\prime}$ are concatenated with the input features ($x_{in}$) and sent to the next feature-dense unit. These learned features from each unit ($x^{\prime}_{1},x^{\prime}_{2},...,x^{\prime}_{i}$) are kept aside and concatenated at the end as the final output of the feature-dense blocks, see Fig. 4 (f). This design has the same number of dense connections as the simple dense blocks but only uses new features for the final output of the block. This allows subsequent layers to focus on new features while utilizing dense connectivity and encouraging better regularization of the cost volume. This design also makes it possible to remove the dropout layer from it. Each feature-dense block follows a similar formulation as in Equation (1), but generates a different output,
$x_{out}=concat\left[x^{\prime}_{0},x^{\prime}_{1},...,x^{\prime}_{i-1}\right],$ (2)
where subscript $i$ is the number of feature-dense units in the block. The number of feature-dense units for each block is shown as a superscript in the U-Net architecture (Fig. 4 (a)). The encoder and the decoder dense blocks have the same number of basic units at the same level. The growth rate $g$ is fixed at $4$ throughout the network.
Figure 6: The visual comparison presents ground truth depth maps for scene 114 from the DTU dataset, showcasing various viewpoints before and after depth filtering. The top row displays the original ground truth depth maps from different viewpoints, while the bottom row illustrates the same maps following filtration. The left half of the figure highlights depth maps with the highest impact (error) from the filtration process, whereas the right half features depth maps with the lowest impact (error).
3.3.4 Transition-down and Transition-up Blocks
The feature maps are downsampled with transition blocks to reduce the 3D spatial resolution (i.e., disparity, height, and width) of the feature maps. Fig. 4 (b) shows the architecture of the transition block. It is a three-layer block, a 1x1-3D convolution followed by a dropout and a 3D max-pooling layer. Similar to the layers inside the dense blocks, the transition block also has a weight-standardized 3D convolution layer. Transition-up layers use a single weight-standardized transpose convolution layer to double the spatial resolution using 3x3x3 kernels. After each transition-up layer, the features from the same level are added together as shown in Fig. 4(a).
3.4 Deformable Feature Extraction Network
Besides the geometric consistency module and dense-CostRegNet, we adopted a new architecture for the Feature Extraction Network (FEN) with weight standardized [55] deformation convolution layers [47, 56]. Deformable layers are known to adjust their sampling locations based on model requirements [47, 56]. This helps extract better features for accelerated learning. We use a multi-level feature pyramid network design for semantically strong feature extraction [48] at all levels.
Deformable convolutional blocks (Deformable-Conv2d) are the building blocks of the FEN as shown in Fig. 5 (a). Each deformable-conv2d block consists of three layers, a weight-standardized conv-2d layer, followed by a group normalization layer, and a ReLU activation. A total of eight such layers are used to extract image features for the three stages, $\frac{HW}{4},\frac{HW}{2}$ and $HW$, as shown in Fig. 5(a). FEN uses a strided deformable-conv2d layer instead of a max-pooling layer to reduce the feature map size at the start of a new stage. This helps to avoid direct feature loss by parameterless max-pooling compression. At the end of each resolution, a stage-out block is applied to obtain the final output features. It consists of four deformable-conv2d layers, as shown in Fig. 5(c). The lowest resolution features, $\frac{HW}{4}$, are the most semantically strong. We upsample the feature maps (using nearest neighbors) to propagate the strong semantic information to the higher-resolution stage. The upsampled features are then added to the features from the downsampling path at the same level. The lateral $1\times 1$ convolution layer matches the number of channels of these two sets of features. The summed output is used as input for the stage-out block, as shown in Fig. 5(a).
Unlike most MVS methods [5, 6, 10, 13, 57, 58] that use batch normalization [59] during training, we use group normalization throughout our network. As observed in [59], batch normalization provides more consistent and stable training with large batch sizes, but it is inconsistent and has a degrading effect on training with smaller batches. MVS methods are restricted to very small batch sizes, often between 141-4, due to large memory requirements by the 3D-regularization network. Thus, we replaced batch normalization with group normalization layers [54] of group size $4$ across the network. Group normalization performs normalization across a number of channels that is independent of the number of examples in a batch [54]. We also implement weight standardization [55] for all layers (including 3D-convolutional layers) in the network. With these modifications, we achieve stable and reproducible training.
3.5 Global Feature Aggregation
Multi-view stereo matching is a one-to-many matching problem, which requires simultaneous consideration of all source views for effective matching. Following TransMVSNet [5], we use the global feature aggregation (GFA) technique just before cost volume creation. GFA aggregates global context information using inter- and intra-image feature interactions [5, 60], as shown in Fig. 5(d). It has been proven improve prediction quality and reduce matching uncertainties, especially for regions with little texture or repetitive patterns.
GFA uses linear attention [61] with multiple heads [34] to estimate attention scores using Query and Key,
$Attention(Q,K,V)\,=\Psi(Q)(\Psi(K^{T})V),$ (3)
where $\Psi(.)=elu(.)+1$, $elu(.)$ represents the activation function of exponential linear unit [5, 62], and V is the value in attention calculation.
The GFA module aggregates global features in two distinct manners: intra-image, in which the Qs and Ks are from the same image (view), and inter-image, when they are from different views. Following TransMVSnet, both the source and reference view features are updated during intra-image GFA, but only the source features are updated for inter-image GFA.
The top of the Fig. 5(d) shows intra-image aggregation. The first row shows the aggregation process for the reference features and the second row shows the aggregation for the source features. The darkness of the arrows indicate the magnitude of attention. The bottom of Fig. 5(d) shows inter-image aggregation where the features from the reference and source images interact to estimate the attention score for global feature aggregation. The aggregated reference and source view features are then used to create the cost volume.
3.6 Cost Function
Most learning-based MVS methods [10, 11, 57] treat depth estimation as a regression problem and use an $L_{1}$ loss between prediction and ground truth. Following AA-RMVSNet [13] and UniMVSNet [6], we treat depth estimation as a classification problem and adopt a cross-entropy loss formulation from AA-RMVSNet [13] (see [6] for relative advantages of regression and classification approaches). The pixelwise depth error $\xi_{d}$ is calculated at each stage,
$\xi_{d}=\mathcal{D}(D^{gt}_{0},D_{0}),$ (4)
where $D^{gt}_{0}$ is the reference ground truth and $D_{0}$ is the reference depth estimate. $\mathcal{D}$ denotes the cross-entropy function modified to produce per-pixel depth error between $D^{gt}_{0}$ and $D_{0}$. We further enhance the one-hot supervision by penalizing each pixel for its inconsistency across different source views. This is implemented using element-wise multiplication ($\odot$) between $\xi_{d}$ and $\xi_{p}$ at each stage. The mean stage loss, $L_{i}$, is calculated as,
$\begin{split}{L_{i}}_{(stage)}&=mean(\xi_{p}\odot\xi_{d})\\ \mathcal{L}_{total}&=\alpha.L_{1}+\beta.L_{2}+\gamma.L_{3}\end{split}$ (5)
where ${L_{i}}_{(stage)}$ is the mean stage loss, $\mathcal{L}_{total}$ is the total loss, and $\alpha,\beta$ and $\gamma$ are the stage-wise weights. This formulation of the cost function with pixel-level inconsistency penalties explicitly forces the model to learn to produce multi-view, geometrically-consistent depth maps.
3.7 Depth Filtering
In analyzing the penalty mask $\xi_{p}$, we observed that certain regions of the scene consistently receive penalties close to the maximum throughout training (see the final $\xi_{p}$ in Fig. 6). This contrasts sharply with other regions, which exhibit a rapid decrease in penalties as training progresses (lighter regions in the same figure). This discrepancy is closely tied to the methods used by MVSNet [15] and R-MVSNet [8] for generating ground truth depth maps.
The DTU dataset [21] provides ground truth point clouds with normal information, which is employed in Screened Poisson Surface Reconstruction (SPSR) [63] to create mesh surfaces. These surfaces are then rendered from each viewpoint to produce depth maps. We suspect that this ground truth generation process introduces geometric inconsistencies into the depth estimates. Consequently, since we use the source view ground truth to calculate per-pixel penalties $\xi_{p}$, this initial error is implicitly fed into the model. This creates a cycle of incorrect feedback for geometric consistency checks, resulting in persistently high penalties for certain regions.
To mitigate the impact of inconsistent depth values, we apply the same GC module (refer to Alg. 1) to filter accurate depth maps, with a minor adjustment. Specifically, we substitute the estimated reference depth map ($D_{0}$) with its ground truth counterpart ($D_{0}^{gt}$) in Alg. 1. The hyperparameters are set to $D_{pixel}$=2, $D_{depth}$=0.25, and MM=8, ensuring that only the most erroneous depth values are filtered out. Fig. 6 displays the original and filtered depth maps from various viewpoints of a sample scene in the DTU dataset. The left half of the figure illustrates viewpoints with the most filtered-out values, while the right half shows viewpoints with the fewest. Although we initially identified this depth map rendering issue in the DTU dataset, the approach is broadly applicable to other datasets employing similar methods for generating ground truth depth maps. For the BlendedMVS dataset, we use $D_{pixel}$=0.5, $D_{depth}$=0.05, and MM=10 in the depth filtering process.
Table 1: Quantitative results on DTU and BlendedMVS. Accuracy (Acc), completeness (comp) and overall are in mmmm. We follow Darmon et al. [64] for BlendedMVS evaluation. Bold and underline represents first and second place, respectively. Error maps are shown in Appendix H.
DTU Dataset BlendedMVS Dataset
Method Acc $\downarrow$ Comp $\downarrow$ Overall $\downarrow$ EPE $\downarrow$ $e_{1}\downarrow$ $e_{3}\downarrow$
Traditional Furu [2] 0.613 0.941 0.777
Tola [4] 0.342 1.190 0.766
Gipuma [1] 0.283 0.873 0.578
COLMAP [3] 0.400 0.664 0.532
Learning-based MVSNet [15] 0.396 0.527 0.462 1.49 21.98 8.32
CasMVSNet [10] 0.325 0.385 0.355 1.43 19.01 9.77
CVP-MVSNet [11] 0.296 0.406 0.351 1.90 19.73 10.24
Vis-MVSNet [57] 0.369 0.361 0.365 1.47 15.14 5.13
AA-RMVSNet [13] 0.376 0.339 0.357
EPP-MVSNet [65] 0.413 0.296 0.355 1.17 12.66 6.20
UniMVSNet [6] 0.352 0.278 0.315
TransMVSNet [5] 0.321 0.289 0.305 0.73 8.32 3.62
GBi-Net [66] 0.315 0.262 0.289
MVSTER [67] 0.350 0.276 0.313
GeoMVSNet [46] 0.331 0.259 0.295
MVSFormer [60] 0.327 0.251 0.289
MVSFormer++ [68] 0.309 0.252 0.2805
GoMVS [69] 0.347 0.227 0.287
GC-MVSNet [19] 0.330 0.260 0.295 0.48 7.48 2.78
GC-MVSNet++ (ours) 0.319 0.246 0.2825 0.407 5.79 2.41
3.8 GC-MVSNet++ Design Insights
The typical supervised MVS pipeline has three interconnected parts: feature extraction, cost volume regularization, and the supervision signal. The feature extraction networks generate initial features used to form the initial cost volume, which is then refined through cost volume regularization to estimate depth. The learning process is driven by the supervision signal provided by the loss function.
Our method begins by enhancing supervision with geometric consistency, explicitly penalizing depth estimates lacking consistency across multiple views. This geometry-aware supervision encourages improved feature extraction with better physical understanding of the 3D scenes and more accurate initial cost volume creation. However, the improved feature extraction and initial cost volume are limited by the rigidity of sampling locations in convolutional layers and the lack of global aggregation. To address this, we first add a global feature aggregation module for intra-image and inter-image feature fusion, and then adopt deformable convolution layers to allow flexible sampling that better accommodates geometric and view variations. (See Table 7 for ablation experiments.)
Inspired by the benefits of flexible feature sampling, we sought a parallel form of flexibility within the cost volume regularization network. Dense network connectivity, as exemplified by DenseNet [52], naturally supports flexible, multi-scale feature reuse with strong gradient flow for optimization [53, 52, 70], thereby enhancing learning representational power. Leveraging this insight, we adopted dense connectivity inspired modules for increased feature reuse, stronger gradient flow, and improved landscape for optimization [71, 70]. While the encoding module (simple dense block) closely follows a traditional DenseNet for its enhanced encoding capabilities [52], we designed a feature-dense decoding module specifically tailored for cost volume refinement. As illustrated in Fig. 4(f), it consists of multiple feature-dense block units, each containing a 3D convolution, group normalization, and ReLU activation. Outputs from these units are concatenated, progressively expanding the receptive field and facilitating accurate estimation across multiple depth planes. This compact yet effective design significantly improves cost volume regularization performance.
Together, these modules form a powerful pipeline comprising a flexible feature extraction network with global feature aggregation, and a robust cost regularization network featuring a densely connected encoder for rich representation learning and a feature-dense decoder with an expanding receptive field to effectively handle multiple depth planes estimates. Empirical validations underscore the efficacy of these components, collectively driving significant improvements in the MVS pipeline.
Table 2: Quantitative results on Tanks and Temples [22]. Bold and underline represents first and second place, respectively.
Intermediate set Advanced set
Method Mean $\uparrow$ Fam. Fra. Hor. Lig. M60 Pan. Pla. Tra. Mean $\uparrow$ Aud. Bal. Cour. Mus. Pal. Tem.
COLMAP [3] 42.14 50.41 22.25 26.63 56.53 44.83 46.97 48.53 42.04 27.24 16.02 25.23 34.70 41.51 18.05 27.94
R-MVSNet [8] 50.55 73.01 54.56 43.42 43.88 46.80 46.69 50.87 45.25 29.55 19.49 31.45 29.99 42.31 22.94 31.10
CasMVSNet [10] 56.84 76.37 58.45 46.26 55.81 56.11 54.06 58.18 49.51 31.12 19.81 38.46 29.10 43.87 27.36 28.11
AA-RMVSNet [13] 61.51 77.77 59.53 51.53 64.02 64.05 59.47 60.85 54.90 33.53 20.96 40.15 32.05 46.01 29.28 32.71
UniMVSNet [6] 64.36 81.20 66.43 53.11 63.46 66.09 64.84 62.23 57.53 38.96 28.33 44.36 39.74 52.89 33.80 34.63
TransMVSNet [5] 63.52 80.92 65.83 56.94 62.54 63.06 60.00 60.20 58.67 37.00 24.84 44.59 34.77 46.49 34.69 36.62
GBi-Net [66] 61.42 79.77 67.69 51.81 61.25 60.37 55.87 60.67 53.89 37.32 29.77 42.12 36.30 47.69 31.11 36.39
MVSTER [67] 60.92 80.21 63.51 52.30 61.38 61.47 58.16 58.98 51.38 37.53 26.68 42.14 35.65 49.37 32.16 39.19
GeoMVSNet [46] 65.89 81.64 67.53 55.78 68.02 65.49 67.19 63.27 58.22 41.52 30.23 46.53 39.98 53.05 35.98 43.34
MVSFormer [60] 66.37 82.06 69.34 60.49 68.61 65.67 64.08 61.23 59.53 40.87 28.22 46.75 39.30 52.88 35.16 42.95
MVSFormer++ [68] 67.03 82.87 68.90 64.21 68.65 65.00 66.43 60.07 60.12 41.70 30.39 45.85 39.35 53.62 35.34 45.66
GoMVS [69] 66.44 82.68 69.23 63.19 63.56 65.13 62.10 58.81 60.80 43.07 35.52 47.15 42.52 52.08 36.34 44.82
GC-MVSNet [19] 62.74 80.87 67.13 53.82 61.05 62.60 59.64 58.68 58.48 38.74 25.37 46.50 36.65 49.97 35.81 38.11
GC-MVSNet++ 66.28 82.72 71.05 58.18 65.40 65.63 64.95 61.47 60.87 42.14 27.74 47.84 40.24 52.32 37.48 47.23
4 Experiments
We evaluate on three datasets of different complexities. DTU [21] is an indoor dataset that contains $128$ scenes with $49$ or $64$ views under $7$ lighting conditions and predefined camera trajectories. We use the same training, validation, and test splits as MVSNet [15]. BlendedMVS [20] is a large-scale synthetic dataset with $113$ indoor and outdoor scenes. It has 106 training scenes and $7$ validation scenes. Tanks and Temples [22] is collected from a more complicated and realistic scene, and contains $8$ intermediate and $6$ advanced scenes. DTU and Tanks & Temples evaluate using point clouds while BlendedMVS evaluates on depth maps.
4.1 Implementation Details
Following general practice [5], we first train and evaluate our model on DTU. Then, we finetune on BlendedMVS to evaluate on Tanks and Temples. For training on DTU, we set the number of input images to $N=5$ and image resolution to $512\times 640$. The depth hypotheses are sampled from $425mm$ to $935mm$ for coarse-to-fine regularization with the number of plane sweeping depth hypotheses for the three stages set to 48, 32, and 8. The corresponding depth interval ratio (DIR) is set as 2.0, 0.8, and 0.4. The model is trained with Adam [72] for 9 epochs with an initial learning rate ($LR_{DTU}$) of 0.001, which decays by a factor of 0.5 once after $8^{th}$ epoch. For the GC-module, we use MM=8 and set the stage-wise thresholds $D_{pixel}$ as 1, 0.5, 0.25 and $D_{depth}$} as 0.01, 0.005, 0.0025. We use $\alpha$$\beta$=1 and $\gamma=2$ for all experiments. We train our model with a batch size of 2 on 8 NVIDIA RTX A6000 GPUs for about 9 hours.
4.2 Experimental Performance
Evaluation on DTU: We generate depth maps with NN=5 at an input resolution of $864\times 1152$. We slightly adjust the depth interval ratio (DIR) to $1.6,0.7,0.3$ to accommodate the resolution change (more on DIR in Appendix C) and use the Fusibile algorithm [1] for depth fusion. Table 1 shows quantitative evaluations, where accuracy is the mean absolute distance in $mm$ from the reconstructed point cloud to the ground truth point cloud, completeness measures the opposite (more details in Appendix E), and overall is the average of these metrics indicate the overall performance of the models. We find that GC-MVSNet++ achieves the second best overall and completeness scores when compared to previous state-of-the-art techniques. Snapshots of the DTU test set point clouds are shown in the Supplementary Materials. We find that our model generates denser and more complete point clouds.
Evaluation on BlendedMVS: Unlike DTU and Tanks and Temples, evaluation on Blended MVS is usually measured as the quality of depth maps, not the quality of point clouds. First, we finetune our model for 12 epochs with NN=5, MM=8, number of depth planes DD=128, at a resolution of $576\times 768$, with one-tenth the learning rate of DTU ($\frac{1}{10}LR_{DTU}$).
Following evalutation process of Darmon et al. [64], we generate Table 1 for a quantitative comparison with other methods using three metrics: Endpoint error (EPE) is the average $L_{1}$ distance between the estimated and the ground truth depth values, and $e_{1}$ and $e_{2}$ are the ratio of number of pixels with $L_{1}$ error larger than 1mm and 3mm, respectively. The significant jump in depth estimates corroborates our intuition that providing explicit geometric cues during training helps the model be mindful about preserving the geometric consistency of a view during inference. The addition of dense-CostRegNet and the use of the depth filtration module led to significant quantitative and qualitative improvements for GC-MVSNet++. error maps and point clouds are shown in Appendices H and I, respectively.).
Figure 7: Precision and recall comparison with other recent methods for Train (intermediate set) and Temple (advanced set) scene on Tanks and Temples benchmark. The error plots are collected from the benchmark website. $\tau$ is the scene-relevant distance threshold. Darker regions indicate larger errors encountered with regard to $\tau$. GC-MVSNet++ shows visual improvements with brighter regions for both metrics.
[…]
4.3 Ablation Study
We conduct ablation studies to evaluate the importance of the various components of GCMVSNet++. We provide detailed comparisons for inference time and memory requirements with MVSFormer++, TransMVSNet, CasMVSNet, and GC-MVSNet in Appendix F of the supplementary material.
$\mathbf{\xi_{p}}$ Range: $\xi_{p}$ is generated using the mask sum (mask_summask\_sum in Alg. 1), and is the sum of penalties accumulated across the $M$ source views during multi-view geometric consistency check. At this stage, its elements take a discrete value between 0 and $M$. Using mask sum as-is leads to a very high penalty per-pixel, and thus a very high loss value, destabilizing the learning process. We control the magnitude of the penalty by controlling the range of the per-pixel penalty.
We explore two different ranges, $[1,2]$ and $[1,3]$. To generate $\xi_{p}\in[1,2]$, we divide the mask sum by $M$ and then add $1$. To generate $\xi_{p}\in[1,3]$, we divide the mask by $\frac{M}{2}$ and then add $1$. Table 5 shows the impact of these two ranges for MM=8. Since $\xi_{p}\in[1,2]$ produces the best results, we use it for all other experiments.
Hyperparameters of GC module The GC module has two types of hyperparameters, global and local. In this section, we investigate the effect of these hyperparameters on our results.
The global hyperparameter $M$ is the number of source views across which the geometric consistency is checked, and is the same for all three stages (coarse, intermediate, and refinement stages). For training on DTU, we vary the value of $M$ while keeping NN=5, i.e. while the MVS method uses only $4$ source views to estimate $D_{0}$, the GC module checks the geometrical consistency of $D_{0}$ across $M$ source views. It is important to note that the first NN-1 out of $M$ source views are exactly the same used by GC-MVSNet++ to estimate $D_{0}$. We always keep $M\geq N-1$.
Table 5 presents a quantitative comparison for different values of $M$ and the number of training iterations required for optimal performance of the model. At MM=N1N-1= 44, i.e. checking geometric consistency across the same number of source views as used by GC-MVSNet++ to estimate $D_{0}$, the model performance significantly improves with a sharp decrease in training iteration requirements, as compared to our baseline GC-MVS-base. As we increase the value of $M$ from $3$ to $10$, the training iterations required by our model further decrease. We find that at $M=8$, which is twice the number of source views used by GC-MVSNet++, it achieves its best performance.
The two local hyperparameters, $D_{pixel}$ and $D_{depth}$, are the stage-wise thresholds applied to generate $PDE_{mask}$ and $RDD_{mask}$ in Alg.1. These are set to smaller values in the later (finer) stages, providing a stricter penalty to geometrically inconsistent pixels at finer resolutions. Table 5 shows the overall performance of GC-MVSNet++ with a range of different $D_{pixel}$ and $D_{depth}$ thresholds. GC-MVSNet++ performance remains fairly consistent and achieves its best performance with $D_{pixel}=1,0.5,0.25$ and $D_{depth}=0.01,0.005,0.0025$. We use these threshold values throughout the paper.
Table 3: Impact of range of $\xi_{p}$ during training on DTU with MM=8, NN=5. Numbers are generated on DTU evaluation set.
[…]
$[1,3]$ 0.331 0.270 0.3005
$[1,2]$ 0.330 0.260 0.295
Table 4: Quantitative results on DTU evaluation set [21]. M is the number of source views used by the GC module for checking the geometric consistency of the reference view depth map. The training iteration requirement of the model decreases as M increases.
[…]
3 0.334 0.274 0.304 14
4 0.343 0.264 0.3035 12
5 0.342 0.271 0.3065 13
6 0.326 0.271 0.298 9
7 0.332 0.270 0.301 10
8 0.330 0.260 0.295 9
9 0.328 0.280 0.304 9
10 0.329 0.268 0.298 10
Table 5: Overall score on the evaluation set of DTU [21] for different values of $D_{depth}$ and $D_{pixel}$. $M$ is fixed at 88. C, I, and R means Coarse, Intermediate, and Refine stages.
[…]
Coarse Inter. Refine Coarse Inter. Refine
0.04 0.03 0.02 4 3 2 0.302
0.03 0.0225 0.015 3 2.25 1.5 0.302
0.02 0.015 0.01 2 1.5 1.0 0.298
0.01 0.008 0.006 1 0.8 0.6 0.303
0.01 0.005 0.0025 1 0.5 0.25 0.295
0.008 0.003 0.002 0.8 0.3 0.2 0.303
0.005 0.002 0.001 0.5 0.2 0.1 0.3015
GC-module as a plug-in: Geometric Consistency module is generic and can be integrated into many different MVS pipelines To demonstrate this, we test it with two very different MVS pipelines, TransMVSNet and CasMVSNet. CasMVSNet treats depth estimation as a regression problem, while TransMVSNet treats it as a classification problem and uses winner-take-all to estimate the final depth map. We purposefully choose different methods to show that the GC module can perform well for both types of formulation. We compare the architectures of GC-MVSNet++ with TransMVSNet and CasMVSNet in discussion section. We also compare it with MVSFormer and MVSFormer++.
Table 7 presents the results, showing the impact of adding the GC module, the deformable feature extraction network (shown as other modifications in the table), and depth filtration in the original method pipeline. To observe the absolute impact of adding these modifications, we do not modify the original pipelines in any other way. We do not include our proposed dense-CostRegNet or Deformable feature extraction network in the original methods as it would enhance their learning capabilities. We focus on the extent of improvement in the original method with the integration of the GC module and depth filtration preprocessing step. We observe in the table that applying only the other modification leads to degradation in performance for both methods. It indicates that the other modification helps in stabilizing the training process and promoting reproducibility but has no significant impact on the performance of the model on its own. We also observe a sharp increase in model performance and a decrease in training iteration requirements after integrating our GC module into the original pipeline. With GC, training the CasMVSNet pipeline requires only $11$ epochs instead of $16$ epochs, while TransMVSNet (with GC module) requires only $8$ epochs instead of $16$ epochs. This corroborates our hypothesis that multi-view geometric consistency significantly reduces training computation because it accelerates the geometric cues learning. Applying the depth filtration preprocessing along with the GC module further improves the performance of both methods. Eliminating the erroneous ground truth pixels from the learning process through filtration provides more consistent geometric cues learning.
Table 6: The impact of GC as a plug-in module on CasMVSNet and TransMVSNet on DTU [21]. Apart from other changes, including a GC module alongside depth filtration (DF) in the original methods boosts the performance. $L_{1}$ and D-FEN indicate $L_{1}$ loss, Focal loss [73], and Deformable Feature Extraction Network, respectively. DF is the depth filtering pre-processing step to remove erroneous ground truth values.
[…]
CasMVSNet $L_{1}$ $\times$ $\times$ $\times$ 0.355 16
$L_{1}$ $\times$ $\checkmark$ $\times$ 0.357 16
$L_{1}$ $\times$ $\times$ $\checkmark$ 0.335 11
$L_{1}$ $\checkmark$ $\times$ $\checkmark$ 0.330 11
TransMVSNet FL $\times$ $\times$ $\times$ 0.305 16
FL $\times$ $\checkmark$ $\times$ 0.322 16
FL $\times$ $\times$ $\checkmark$ 0.303 8
FL $\checkmark$ $\times$ $\checkmark$ 0.297 8
Table 7: Stages of performance improvement of GC-MVSNet++ with different modifications on DTU [21]. It uses Cross-entropy loss [13] for training. DF, D-FEN, GC, GFA, and DenseCostReg indicate depth filtering, Deformable-Feature Extraction Network, Geometric Consistency module, Global Feature Aggregation and Dense Cost Regularization network, respectively.
[…]
GC-MVS-Base $\times$ $\times$ $\times$ $\times$ $\times$ 0.332
GC-MVS-Base $\times$ $\checkmark$ $\times$ $\times$ $\times$ 0.334
GC-MVS-Base $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ 0.328
GC-MVS-Base $\times$ $\times$ $\checkmark$ $\times$ $\times$ 0.298
GC-MVS-Base $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ 0.295
GC-MVS-Base $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 0.291
GC-MVSNet++ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.2825
Stages of GC-MVSNet++: Table 7 shows different stages of development of GC-MVSNet++. GC-MVS-base uses the TransMVSNet pipeline (without the feature matching transformer and adaptive receptive field modules) with cross-entropy loss and performs much worse than original TransMVSNet which uses focal loss. Including the global feature aggregation module with the basic feature extraction network performs roughly similar. But including D-FEN (Deformable-Feature Extraction Network) modifications slightly improves the overall performance indicating better feature quality with deformable convolution for global feature aggregation. However, D-FEN does not reduce the training iteration requirements on its own, see Table. 7. Only after applying the GC module, independently and with D-FEN, we see a significant reduction in training iteration requirements and a significant improvement in the overall accuracy metric. This clearly shows the significance of multi-view multi-scale geometric consistency checks in the GC-MVSNet++ pipeline. Including the DF preprocessing step further aids the GC module by eliminating erroneous ground truth values from the training loop and provides more robust geometry cues across multiple source views. Finally, adding dense-CostRegNet with all other modifications improves the cost regularization process with its advanced architecture and dense identity connection, leading to a significant improvement.
5 Discussion
We provide a comparison of our model with recent state-of-the-art models like MVSFormer, MVSFormer++, TransMVSNet and CasMVSNet.
GC-MVSNet++ vs. MVSFormer and MVSFormer++: MVSFormer focuses on using state-of-the-art pre-trained models for feature extraction and extensively use transformers in all its components. MVSFormer++ further adds multiple transformer modules in the MVS pipeline to leverage global feature aggregation in cost volume regularization. GC-MVSNet++ uses only one global feature aggregation module after feature extraction for global context enhancement before cost volume creation. It focuses on enforcing geometry-based learning across multiple views and on effective utilization of dense connections to enhance the cost regularization network.
GC-MVSNet++ vs. TransMVSNet: TransMVSNet uses regular 2D convolution-based FPN (with batch-norm) for feature extraction and employs adaptive receptive field (ARF) modules with deformable layers after feature extraction. It trains using focal loss [73]. GC-MVSNet++ replaces the combination of regular FPN and ARF modules with weight-standardized deformable FPN (with group-norm) for feature extraction. It also uses a novel densely connected CostRegNet for cost volume regularization. It trains with a combination of cross-entropy loss and geometric consistency penalty for accelerated learning.
GC-MVSNet++ vs. CasMVSNet: CasMVSNet [10] proposes a coarse-to-fine regularization technique. It uses a regular 2D convolution-based FPN for feature extraction, generates variance-based cost volume, and employs depth regression to estimate $D_{0}$. The only similarity with our model is that we also use coarse-to-fine regularization.
6 Conclusion
In this paper, we introduce GC-MVSNet++, an enhanced learning-based MVS pipeline that explicitly models the geometric consistency of reference depth maps across multiple source views during training. Our approach incorporates a novel dense-CostRegNet with two distinct modules, simple dense-module which is effective for encoding and feature-dense-module, which is designed to leverage dense connection for effective decoding and precise estimation of reference view depth maps. Ours is the first few method to integrate multi-view, multi-scale geometric consistency checks into the training process. We demonstrate that the GC module is versatile and can be integrated with other MVS methods to enhance their learning. Through extensive experiments and ablation studies, we highlight the advantages of GC-MVSNet++. This work shows how to blend traditional geometric techniques and modern machine learning methods, to achieve more accurate and reliable 3D reconstructions from multiple images.
Acknowledgement: This work was supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by the government of the Republic of Korea [25ZC1110, The research of the basic media $\cdot$ contents technologies].